RhondaSommer (rhondasommer):

What is the average kinetic energy of 1 mole of a gas at -32 degrees Celsius? (R = 8.314 J/K-mol) \[3.01~x~10^3 J\]\[-3.99~x~10^2 J\]\[6.01~x~10^3 J\]\[4.32~x~10^3 J\]

RhondaSommer (rhondasommer):

I guessed and got it wrong :/

OpenStudy (math&ing001):

Use the formula : \[KE_{avg}=\frac{3}{2}RT \] Don't forget to turn the temperature to kelvin ;]

OpenStudy (math&ing001):

@RhondaSommer you wanna try again ?

RhondaSommer (rhondasommer):

im sooo confused by that formula

OpenStudy (math&ing001):

KE stands for Kinetic Energie (what you're looking for :P) T stands for temprature, and you already got R (the universal gas constant) Still anything confusing you ?

RhondaSommer (rhondasommer):

KE = 3/2 * n * R T?

RhondaSommer (rhondasommer):

KE = 3/2 * 1 * 8.314 * (-32 + 273.15)

OpenStudy (math&ing001):

Why "n" ? That's kinetic energie per mol..

OpenStudy (math&ing001):

Yep, looks correct !